# Diffusion Dataset Condensation: Training Your Diffusion Model Faster with Less Data

### v1

@@ 3.3 D2C Training Process
Table 1: Comparison of gFID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet 256$\times$256. We use CFG=1.5 for evaluation. D2C surpasses other methods at all settings.
@@ 4 Experiments
Figure 6: Left: Ablation studies of the interval value $k$ in interval sampling at different data budgets. With a small interval value, training can experience significant acceleration during the early iterations. However, due to the limited diversity of the selected dataset, performance tends to be overtaken by random sampling in later iterations. As the interval value increases, the final gFID-10K scores initially improve and then decline, suggesting an optimal interval value exists. Notably, this optimal value is 96 for a 10K data budget and 16 for a 50K data budget, and the ratio 96/16 closely approximates the ratio of the data budgets, 50K/10K. Right:. Ablation study of DC-Embedding under a 10K data budget. “Only Class” stands for the baseline approach of injecting class embeddings. This figure demonstrates that both the text embedding branch and the original class embedding branch (see Fig. 5) are critical for the effective training of diffusion models.
@@
Evaluation and baselines. We train models from scratch on the collected subset and evaluate them using gFID [38], sFID, Inception Score [39], Precision (Prec.), and Recall (Rec.), adhering to standard evaluation protocols [16, 11, 13]. We compare our method against REPA [12], and various data condensation and selection baselines, including SRe2[24], Herding, K-Center, and random sampling, using SiT and DiT architectures [13, 11]. Further details regarding evaluation metrics and baseline methods can be found in Appendix G and H.
@@ 4.1 Main Result
Training Performance and Speed. We evaluate D2C on SiT-XL/2 using 10K and 50K data budgets, comparing its performance against REPA and a vanilla SiT model trained on the full ImageNet dataset (a 1.28M data budget), as well as random selection with 10K and 50K data budgets. As illustrated in Fig. 1 (b), our method achieves a gFID-50K of 4.23 at only 40K iterations with 10K training data and without CFG. In contrast, REPA requires 4 million steps and the vanilla SiT model needs over 7 million steps to reach comparable performance, representing an acceleration of over 100$\times$ and 233$\times$, respectively. Under a 4% data budget (50K) with CFG set to 1.5, our method achieves an FID of 2.78 at 180K steps, further demonstrating significant data and compute efficiency (Fig. 1 (c)). Moreover, Fig. 4 presents a visual comparison between random selection and our D2C at 10K and 50K data sizes, where images are generated from the same noise and class label. Our method demonstrates superior visual quality compared to the baseline and generates higher-quality images, even during the early iterations of training.
Comparison on ImageNet 256$\times$256. We compare D2C with random sampling, Herding [40], K-Center [30], and SRe2[24] under various data budgets and backbones. As illustrated in Table 1, D2C consistently achieves the lowest FID across all settings. For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a gFID-50K of 4.20 on DiT-L/2 and 3.98 on SiT. These results demonstrate the superiority of our approach over existing methods. Notably, SRe2L, which performs well in classification task, fails on this generative task (see Table 3) due to its focus on category-discriminative features. Similarly, geometry-based methods like Herding and K-Center, along with random sampling, prove inadequate for achieving efficient and high-performing training of diffusion models.
Comparison on ImageNet 512$\times$512. As shown in Table 2, D2C achieves a gFID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet 512$\times$512 settings. On SiT-L/2, similar improvements are observed. These results demonstrate that D2C generalizes well to higher resolutions.
@@ 4.2 Ablation Study
Ablation on Select Phase. We investigate the impact of the interval value $k$ in the Select phase, as shown in Fig. 6 (Left) and Fig. 7. Using a small value accelerates early training by prioritizing min-loss samples, which are simpler and easier to learn. However, the limited diversity of such samples leads to degraded performance in later stages, eventually being overtaken by settings with moderate interval values. In contrast, large intervals or random selection introduce excessive max-loss or uncurated samples, destabilizing training (Fig. 3). As $k$ increases, we observe that gFID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability. Empirically, the best results are achieved with an interval of 96 for the 10K budget and 16 for 50K, approximately following the ratio of data budgets (50K/10K).
@@
Ablation on Attach Phase. We conduct ablation studies on the Attach phase from two perspectives. First, as shown in Fig. 6 (Right), we examine the effectiveness of dual conditional embedding under a 10K data budget. Using only class or text embedding results in degraded performance compared to their combination, indicating the necessity of both components to capture category semantics and text embeddings. Notably, text embedding alone performs better than class embedding, suggesting that textual descriptions provide richer semantic information for generative tasks. Second, as shown in Table 4, we analyze the impact of the semantic and visual information injection modules. The gFID-10K increases to 37.07 when both branches are removed and to 9.01 when only the visual injection module is excluded. Notably, the combination of both modules yields the best FID of 7.62. These results underscore the crucial role of injecting both semantic and visual information in accelerating diffusion model training.
@@ Appendix E More Discussions about Select
Figure 8: Left: gFID-10K across training steps under different interval values $k$ for a 50K data budget. Moderate intervals (e.g., $k=16$) achieve superior performance by balancing learnability and diversity. Right: Distributional discrepancy (gFID-10K) between ranked training subsets and the validation set. Both extremely low and high diffusion difficulty scores lead to higher FID, while mid-range segments show better alignment.
@@
The right panel in Fig. 8 presents the gFID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset. We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using gFiD. Interestingly, we observe a clear U-shaped curve: subsets consisting of extremely low or high difficulty samples exhibit significantly worse distributional alignment, while those centered around moderate difficulty levels show substantially lower FID scores. This result aligns well with our hypothesis that very easy samples (e.g., simple textures, clean backgrounds) and extremely hard samples (e.g., ambiguous, noisy structures) both fail to reflect the global data distribution.
@@ Appendix G Evaluation Details
  • gFID [38] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [44].
@@ Appendix I Visualization of SRe2L in Generative Tasks
As shown in Table 5, D2C consistently outperforms random sampling under a strict 10K (0.8%) data budget across both DiT-L/2 and SiT-L/2 backbones. At 300k iterations, our method achieves a gFID of 5.8 on DiT-L/2 and 4.22 on SiT-L/2, showing strong improvements. Visual samples in Figure 11 further confirm the high fidelity and diversity of generations at 512$\times$512 resolution, demonstrating that D2C generalizes effectively to high-resolution settings.

### v2

@@ 3.3 D2C Training Process
Table 1: Comparison of FID-50K across various dataset condensation methods and data budgets using DiT-L/2 and SiT-L/2 on ImageNet 256$\times$256. We use CFG=1.5 for evaluation. D2C surpasses other methods at all settings.
@@ 4 Experiments
Figure 6: Left: Ablation studies of the interval value $k$ in interval sampling at different data budgets. With a small interval value, training can experience significant acceleration during the early iterations. However, due to the limited diversity of the selected dataset, performance tends to be overtaken by random sampling in later iterations. As the interval value increases, the final FID-10K scores initially improve and then decline, suggesting an optimal interval value exists. Notably, this optimal value is 96 for a 10K data budget and 16 for a 50K data budget, and the ratio 96/16 closely approximates the ratio of the data budgets, 50K/10K. Right:. Ablation study of DC-Embedding under a 10K data budget. “Only Class” stands for the baseline approach of injecting class embeddings. This figure demonstrates that both the text embedding branch and the original class embedding branch (see Fig. 5) are critical for the effective training of diffusion models.
@@
Evaluation and baselines. We train models from scratch on the collected subset and evaluate them using FID [38], sFID, Inception Score [39], Precision (Prec.), and Recall (Rec.), adhering to standard evaluation protocols [16, 11, 13]. We compare our method against REPA [12], and various data condensation and selection baselines, including SRe2[24], Herding, K-Center, and random sampling, using SiT and DiT architectures [13, 11]. Further details regarding evaluation metrics and baseline methods can be found in Appendix G and H.
@@ 4.1 Main Result
Training Performance and Speed. We evaluate D2C on SiT-XL/2 using 10K and 50K data budgets, comparing its performance against REPA and a vanilla SiT model trained on the full ImageNet dataset (a 1.28M data budget), as well as random selection with 10K and 50K data budgets. As illustrated in Fig. 1 (b), our method achieves a FID-50K of 4.23 at only 40K iterations with 10K training data and without CFG. In contrast, REPA requires 4 million steps and the vanilla SiT model needs over 7 million steps to reach comparable performance, representing an acceleration of over 100$\times$ and 233$\times$, respectively. Under a 4% data budget (50K) with CFG set to 1.5, our method achieves an FID of 2.78 at 180K steps, further demonstrating significant data and compute efficiency (Fig. 1 (c)). Moreover, Fig. 4 presents a visual comparison between random selection and our D2C at 10K and 50K data sizes, where images are generated from the same noise and class label. Our method demonstrates superior visual quality compared to the baseline and generates higher-quality images, even during the early iterations of training.
Comparison on ImageNet 256$\times$256. We compare D2C with random sampling, Herding [40], K-Center [30], and SRe2[24] under various data budgets and backbones. As illustrated in Table 1, D2C consistently achieves the lowest FID across all settings. For instance, using only 0.8% of the data and 100K iterations with early stopping, our method achieves a FID-50K of 4.20 on DiT-L/2 and 3.98 on SiT. These results demonstrate the superiority of our approach over existing methods. Notably, SRe2L, which performs well in classification task, fails on this generative task (see Table 3) due to its focus on category-discriminative features. Similarly, geometry-based methods like Herding and K-Center, along with random sampling, prove inadequate for achieving efficient and high-performing training of diffusion models.
Comparison on ImageNet 512$\times$512. As shown in Table 2, D2C achieves a FID of 5.8 on DiT-L/2, a significant improvement over the 17.1 achieved by random sampling at 300k iterations under the ImageNet 512$\times$512 settings. On SiT-L/2, similar improvements are observed. These results demonstrate that D2C generalizes well to higher resolutions.
@@ 4.2 Ablation Study
Ablation on Select Phase. We investigate the impact of the interval value $k$ in the Select phase, as shown in Fig. 6 (Left) and Fig. 7. Using a small value accelerates early training by prioritizing min-loss samples, which are simpler and easier to learn. However, the limited diversity of such samples leads to degraded performance in later stages, eventually being overtaken by settings with moderate interval values. In contrast, large intervals or random selection introduce excessive max-loss or uncurated samples, destabilizing training (Fig. 3). As $k$ increases, we observe that FID-10K first decreases and then worsens, revealing an optimal trade-off between diversity and learnability. Empirically, the best results are achieved with an interval of 96 for the 10K budget and 16 for 50K, approximately following the ratio of data budgets (50K/10K).
@@
Ablation on Attach Phase. We conduct ablation studies on the Attach phase from two perspectives. First, as shown in Fig. 6 (Right), we examine the effectiveness of dual conditional embedding under a 10K data budget. Using only class or text embedding results in degraded performance compared to their combination, indicating the necessity of both components to capture category semantics and text embeddings. Notably, text embedding alone performs better than class embedding, suggesting that textual descriptions provide richer semantic information for generative tasks. Second, as shown in Table 4, we analyze the impact of the semantic and visual information injection modules. The FID-10K increases to 37.07 when both branches are removed and to 9.01 when only the visual injection module is excluded. Notably, the combination of both modules yields the best FID of 7.62. These results underscore the crucial role of injecting both semantic and visual information in accelerating diffusion model training.
@@ Appendix E More Discussions about Select
Figure 8: Left: FID-10K across training steps under different interval values $k$ for a 50K data budget. Moderate intervals (e.g., $k=16$) achieve superior performance by balancing learnability and diversity. Right: Distributional discrepancy (FID-10K) between ranked training subsets and the validation set. Both extremely low and high diffusion difficulty scores lead to higher FID, while mid-range segments show better alignment.
@@
The right panel in Fig. 8 presents the FID-10K scores of subsets sampled from different portions of the difficulty-ranked dataset. We partition the training set into consecutive 10K segments ordered by the diffusion difficulty score (e.g., the first 10K samples with lowest scores as “Min”, followed by 10–20K, 20–30K, and so on), and measure each segment’s discrepancy from the full validation distribution using FID. Interestingly, we observe a clear U-shaped curve: subsets consisting of extremely low or high difficulty samples exhibit significantly worse distributional alignment, while those centered around moderate difficulty levels show substantially lower FID scores. This result aligns well with our hypothesis that very easy samples (e.g., simple textures, clean backgrounds) and extremely hard samples (e.g., ambiguous, noisy structures) both fail to reflect the global data distribution.
@@ Appendix G Evaluation Details
  • FID [38] computes the Fréchet distance between the feature distributions of real and generated images. Features are extracted using the Inception-v3 network [44].
@@ Appendix I Visualization of SRe2L in Generative Tasks
As shown in Table 5, D2C consistently outperforms random sampling under a strict 10K (0.8%) data budget across both DiT-L/2 and SiT-L/2 backbones. At 300k iterations, our method achieves a FID of 5.8 on DiT-L/2 and 4.22 on SiT-L/2, showing strong improvements. Visual samples in Figure 11 further confirm the high fidelity and diversity of generations at 512$\times$512 resolution, demonstrating that D2C generalizes effectively to high-resolution settings.